OpenStudy (anonymous):

Solve the following system. x - y = 10 x + y = 8 Answer (9, -1) (9, 1) (-1, 9) (-9, 1)

OpenStudy (anonymous):

Check out your last question I showed you how to solve these types of questions

OpenStudy (anonymous):

(9,-1)

OpenStudy (anonymous):

x = 10 + y so 10 + y + y = 8 2y/2 = -2/2 y = -1 therefore x - 1 = 8 x = 9 therefore, (9,-1) as (x,y) These questions are easy I hope you can follow me so you dont need help anymore with them